Jason invested $12,000 in two funds. One of the funds returned 6% interest, and the other returned 8% interest after one year. If the total interest for the year was $880, how much did he invest in each fund?
Answer by sachi(548)

let investment in one of the fund is $x returned 6% interest, and investment in the other the fund is $(12000-x) returned 8% interest after one year.
so the total interest =6*x/100+8*(12000-x)/100=880 (as per question)
or 6x+8(12000-x)=88000
or -2x=88000-96000=-8000
or x=4000
so 1st investment is 4000 at 6% & other is 8000 at 8%